OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. D(5, -2) m = 2/5

OpenStudy (anonymous):

First, you should figure out the equation of the line in slope-intercept form. You already know the slope m=2/5 Now find the y-intercept by using the point. Equation for a line in slope-intercept form is y=mx+b. Substitute 5 for x and -2 for y. -2 = (2/5)(5) + b Solve for b: b=-4 So the equation of your line is y = (2/5)x -4 You don't want any fractions, though, because you want the equation in general form. Multiply both sides by 5. 5y=2x-20 Put the x-term and y-term on the same side of the equation by subtracting 2x from each side. -2x + 5y = -20 Alternatively, you could have 2x - 5y = 20